OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3)?

OpenStudy (mww):

This line being parallel has the same gradient. So m = 2. Then we use the 1 point formula with the point (5,3) y - y1 = m(x - x1) y - 3 = 2(x - 5) --> y - 3 = 2x - 10 so y =2x -7.

OpenStudy (anonymous):

thank you so much!